OpenStudy (anonymous):

m^2/n^2is equivalent to _____. a) mn^-2 b) m^-2n^-2 c) m^2n^-2 d) m^-2n^2

OpenStudy (lgbasallote):

hint!!! \[\LARGE \frac{1}{a^n} = a^{-n}\] hope that helps :)

OpenStudy (anonymous):

yes, but this equals the denominator, so what is the numerator? :S

OpenStudy (lgbasallote):

the numerator stays the same ;)

OpenStudy (lgbasallote):

but if you wanna know.. \[\LARGE \frac{1}{b^{-n}} = b^n\]

OpenStudy (anonymous):

C is the correct anwer..

OpenStudy (anonymous):

As: \[1/b^2 = b^{-2}\] therefore, \[m^2/n^2 = \]\[m^2 \times 1/n^2 = m^2 \times n^{-2}\]